Annalyn is 3 years older than Alvin. In 7 years, Alvin's age will be 6/7 of Annalyn's age. How old are they now?
Answer by lwsshak3(11628)

Annalyn is 3 years older than Alvin. In 7 years, Alvin's age will be 6/7 of Annalyn's age. How old are they now?**
let x= Alvin's age now
x+3= Annalyn's age now
In 7 years,
x+7= Alvin's age
x+3+7=Annalyn's age
..
(x+7)/(x+10)=6/7
7x+49=6x+60
x=11
x+3=14
ans:
Alvin's age now=11
Annalyn's age now=14